Let Q(x, y) be the statement “x + y = x − y.” If the domain for both variables consists of all integers, what are the truth values? 


1
Expert's answer
2022-04-17T23:13:07-0400

So, 0 is an integer. If we put make 0+0 = 0.

Then, 0-0 = 0, so 0+0 = 0-0

it means that x = 0, y = 0